OpenStudy (anonymous):

Find the area of an equilateral triangle (regular 3-gon) with the given measurement. 6-inch apothem A = sq. in.

OpenStudy (anonymous):

area=A^2*N*tan(pi/N) A= length of apothem=6 N=no. of sides=3 pi/N=180/3=60 area=6^2*3*tan(60) =36*3*sqrt(3) =108sqrt(3)

OpenStudy (anonymous):

ap=6, area A=bh/2, base b=2ap*sqrt3, height h=3ap A=bh/2 A=(2apsqrt3)(3ap)/2 A=3(ap)^2 sqrt3 A=3*36*sqrt3=187.06 square inch